# Double strangeness pentaquarks and other exotic hadrons in the $\Xi_{b}\to\Xi\ J/\Psi\ \phi$ decay

### v1

@@ Abstract
We study the possibility that four $\Xi$ resonances $(\Xi(1620)$, $\Xi(1690)$, $\Xi(1820)$, $\Xi(1950))$ could correspond to pentaquark states, in the form of meson-baryon bound systems. We also explore the possible existence of doubly strange pentaquarks with hidden charm $(P_{css})$ and find two candidates structured in a similar form, at energies of $4493~{}{\rm MeV}$ and $4630~{}{\rm MeV}$. The meson-baryon interaction is built from t-channel meson exchange processes which are evaluated using effective Lagrangians. Moreover we analyse the $\Xi_{b}\to\Xi~{}J/\psi~{}\phi$ decay process, which permits exploring the existence of the heavy double strange pentaquark, as well as other exotic hadrons, in the three different two-body invariant mass spectra of the emitted particles. In the $J/\psi\phi$ mass spectrum, we analyse the nature of the $X(4140)$ and $X(4160)$ resonances. In the $J/\psi\Xi$ invariant mass spectrum, we study the signal produced by the doubly strange pentaquark, where we conclude that it has a good chance to be detected in this reaction if its mass is around $4580-4680~{}{\rm MeV}$. Finally, in the $\phi\Xi$ spectrum we study the likelihood to detect the $\Xi(2500)$ state.
@@ I Introduction
One example of the success of this type of models is provided by the $\Lambda(1405)$ resonance, for which the quark models systematically over-predicted its mass. Instead, dedicated studies of the meson-baryon interaction obtained from chiral effective Lagrangians predict the $\Lambda(1405)$ to be a $\bar{K}N$ molecule Dalitz:1960du ; Kaiser:1995eg ; Oset:1997it , what actually allows to interpret the $\Lambda(1405)$ as the first observed pentaquark. Furthermore, the chiral models with unitarization in coupled channels have predicted its double-pole nature Oller:2000fj ; Jido:2003cb , which can be seen from comparing different experimental line shapes Magas:2005vu , and which is now commonly accepted in the field and appears in the PDG  PDG . After the successful interpretation of the $\Lambda(1405)$, many groups tried to extend this kind of models to other spin, isospin and flavour sectors, finding a more natural way to explain different states, like the $f_{0}(500)$ and the $a_{0}(980)$, which are described as meson-meson quasi-bound states Oller:1997ti ; Oller:1998hw ; Pelaez:2015qba .
@@
More recent experimental results, like the discovery at LHCb LHCb:2015yax ; LHCb:2019kea of four excited resonances of the nucleon ($P_{c}(4312)$, $P_{c}(4380)$, $P_{c}(4440)$ and $P_{c}(4457))$, seen in the invariant mass distribution of $J/\psi p$ pairs from the $\Lambda_{b}$ decay, and the more recent report of LHCb M.Wang:Psc which, analysing the $\Xi_{b}$ decay, finds evidence of the existence of a pentaquark with strangeness $(P_{cs}(4459))$ on the invariant mass distribution of $J/\psi\Lambda$ pairs, clearly establish the need for including a $c\bar{c}$ pair excitation in order to reproduce the high masses of these states.
@@
Baryons with hidden charm, structured as a meson-baryon molecule, had been predicted in some earlier works Wu:2010vk ; Hofmann:2005sw ; Yuan:2012wz ; Xiao:2013yca ; Garcia-Recio:2013gaa , prior to their discovery, but the existence of a hidden charm pentaquark with double strangeness was not addressed until recently. Adopting a one-boson-exchange model to derive effective interaction potentials among the hadrons, the work of Wang:2020bjt found a pair of slightly bound meson-baryon molecules with double strangeness, coupling strongly to $\Xi_{c}^{\prime}\bar{D}_{s}^{*}$ and $\Xi_{c}^{*}\bar{D}_{s}^{*}$ states respectively, albeit with the use of somewhat hard and unrealistic cut-off values of around 2 GeV. In contrast, the works based on a t-channel vector-meson exchange interaction between mesons and baryons in the hidden charm sector did not find double strangeness pentaquark molecules Wu:2010vk . However, a recent study pointed out that, in spite of the fact that these models predict weakly attractive or even repulsive meson-baryon interactions, a strong coupled-channel effect does indeed produce enough attraction to generate bound pentaquarks Marse-Valera:2022khy ; Magas:2024biu with strangeness $S=-2$, $P_{css}$. The prediction of such pentaquarks has been corroborated in Ref. Roca:2024nsi , where a similar t-channel vector-meson interaction was employed, but deriving the baryon-baryon-vector meson vertices from the spin and flavour wave functions rather than from SU(4) arguments as in Marse-Valera:2022khy . We note that $P_{css}$ pentaquarks have also been predicted within quark-model approaches Anisovich:2015zqa ; Ortega:2022uyu or employing sum rule techniques Azizi:2021pbh .
@@
In the recent work of Ref. Oset:2024fbk the authors have proposed the $\Xi_{b}^{0}\to\eta\eta_{c}\Xi^{0}$ and $\Omega_{b}^{-}\to K^{-}\eta_{c}\Xi^{0}$ decay processes to look for the $P_{css}$ with $I(J^{P})=\frac{1}{2}(\frac{1}{2}^{-})$ generated via the pseudoscalar-baryon interaction. The enhanced intensity of the $\Omega_{b}^{-}$ decay process establishes it as having a better chance for observation according to their study, but it is also much more difficult experimentally.
@@
This paper has a twofold purpose. On the one hand, we present the details of the formalism that produces a dynamically generated $P_{css}$’s, which were omitted in a previously published short letter Marse-Valera:2022khy , while showing at the same time the prediction of the model for the lighter energy range of the $S=-2$ sector (including some other resonances generation). This sector has experienced a renewed interest due to a recent data from Belle experiment Belle:2018lws , studied, for example, in Nishibuchi:2023acl ; Feijoo:2023wua , and even more recent from the femtoscopy analysis from the ALICE collaboration ALICE:2023wjz , which are analyzed in particular in Ref. Sarti:2023wlg .
@@
On the other hand, we propose a decay process, $\Xi_{b}\to\Xi~{}J/\psi~{}\phi$, in which the predicted pentaquark can be seen in $J/\psi\Xi$ spectrum. As we will discuss in section IV.1 this reaction is interesting, because it allows to detect and study some other exotic and not only hadrons too, namely a family of $X$ resonances from $X(4140)$ till $X(4700)$ and $\Xi(2500)$.
@@ II.1 Meson-Baryon interaction
The model of the meson-baryon interaction used in this work is based on the tree-level diagrams of Fig. 1. We will only consider the s-wave amplitude, for which the most important contribution comes from the t-channel term (Fig. 1(a)). The s- and u- channel terms (Fig. 1(b) and (c)) contribute mostly to the p-wave amplitude and they may have effects at higher energies. In Ref. Oller:2000fj the contribution from these terms in the $S=-1$ sector were calculated, and were found to be around $20\%$ of that of the dominant t-channel around $200~{}{\rm MeV}$ above the threshold. We can expect that in the $S=-2$ sector studied here, these terms will contribute even less, as the intermediate baryon is more massive, and therefore they will be neglected.
Figure 1: Leading order tree-level diagrams contributing to the meson-baryon interaction. Baryons and mesons are depicted by solid and dashed lines, respectively.
To describe the interaction we employ effective Lagrangians that describe the couplings of the vector meson to pseudoscalar mesons $(VPP)$ and baryons $(VBB)$, which are obtained using the hidden gauge formalism and assuming $SU(4)$ symmetry Hofmann:2005sw :
@@
Using the $VPP$ and $VBB$ vertices in the t-exchange diagram of Fig. 1 (a) one can obtain the s-wave interaction kernel $V_{ij}$ Hofmann:2005sw for the pseudoscalar meson baryon $(PB)$ interaction:
@@ II.2 The $I=1/2~{}S=-2$ sector
In the $I=1/2~{}S=-2$ sector we have $9$ possible pseudoscalar-meson baryon (PB) channels, namely $\pi\Xi(1456)$, $\bar{K}\Lambda(1611)$, $\bar{K}\Sigma(1689)$, $\eta\Xi(1866)$, $\eta^{\prime}\Xi(2276)$, $\eta_{c}\Xi(4302)$, $\bar{D}_{s}\Xi_{c}(4437)$, $\bar{D}_{s}\Xi_{c}^{\prime}(4545)$, $\bar{D}\Omega(4565)$, where the values in parenthesis denote their thresholds. As can be seen, the channels with hidden charm are about $2~{}{\rm GeV}$ more massive compared with the lighter channels. Therefore we can expect these two regions to be essentially independent, making only a rather small effect on each other. The values of the $C_{ij}$ coefficients for the 9-channels are displayed in Table 1.
@@
$\eta^{\prime}\Xi$ 00 00 $\frac{1}{\sqrt{8}}\kappa_{c}$ $-\frac{1}{\sqrt{6}}\kappa_{c}$ $\frac{1}{\sqrt{3}}\kappa_{c}$
@@
In the vector meson baryon (VB) sector, the allowed channels are $\rho\Xi(2089)$, $\bar{K}^{*}\Lambda(2010)$, $\bar{K}^{*}\Sigma(2087)$, $\omega\Xi(2101)$, $\phi\Xi(2338)$, $J/\psi\Xi(4415)$, $\bar{D}^{*}_{s}\Xi_{c}(4581)$, $\bar{D}^{*}_{s}\Xi_{c}^{\prime}(4689)$, $\bar{D}^{*}\Omega(4706)$, where we can again separate them in to groups, one with the light channels and the other with the heavy ones. In this case we can obtain the $C_{ij}$ coefficients directly from those in Table 1 using the following correspondences,
$\begin{matrix}\pi\rightarrow\rho,&K\rightarrow K^{\ast},&\bar{K}\rightarrow\bar{K}^{\ast},&D\rightarrow D^{\ast},&\bar{D}\rightarrow\bar{D}^{\ast}\end{matrix}$
$\begin{matrix}\eta_{c}\rightarrow J/\psi,&\frac{1}{\sqrt{3}}\eta+\sqrt{\frac{2}{3}}\eta^{\prime}\rightarrow\omega,&-\sqrt{\frac{2}{3}}\eta+\frac{1}{\sqrt{3}}\eta^{\prime}\rightarrow\phi\ ,\end{matrix}$ (9)
@@
$\omega\Xi$ 00 00 00 $\kappa_{c}$ $\frac{1}{\sqrt{3}}\kappa_{c}$ $\frac{1}{\sqrt{6}}\kappa_{c}$
$\phi\Xi$ 00 00 $\frac{1}{\sqrt{8}}\kappa_{c}$ $-\frac{1}{\sqrt{6}}\kappa_{c}$ $\frac{1}{\sqrt{3}}\kappa_{c}$
@@
Note that for the light $PB$ sector, due to the fact that all $C_{5j}$ coefficients with $j\leq 5$ are zero, we formally have only $4$ coupled channels, while this is not the case for the light $VB$ sector which remains with $5$ coupled channels in this case.
@@ II.3 Dynamically generated resonances
The sought resonances are dynamically generated as poles of the scattering amplitude $T_{ij}$, which is unitarized via the on-shell Bethe-Salpeter equation in coupled channels. This corresponds to performing the resumation of loop diagrams to an infinite order (see Fig. 2),
@@
The masses $M_{l}$ and $m_{l}$ correspond to those of the baryon and meson of the $l$ channel, respectively, $P=p+k=(\sqrt{s},0)$ is the total four momenta in the c.m. frame, and $q$ denotes the four momentum in the intermediate loop. We factorize the $V$ and $T$ matrices on-shell out of the integrals in Eq. (10) Montana:2017kjw , hence obtaining the following algebraic matrix expression for the scattering amplitude,
@@
Note that the sum over the polarization vector in the case of the vector meson baryon interaction, eq. (8), gives Montana:2017kjw
$\sum_{pol}\epsilon_{i}\epsilon_{j}=\delta_{ij}+\frac{q_{i}q_{j}}{M_{V}^{2}},$ (13)
where, consistently with our model, we neglect the $q^{2}/M_{V}^{2}$ term, thus all $\vec{\epsilon}_{i}\vec{\epsilon}_{j}$ factorize out from all terms in the Bethe-Salpeter equation.
@@ II.4 Loop function
were $q_{l}$ is the on-shell three-momentum of the meson in the loop, and we have introduced a subtraction constant, $a(\mu)$, at the regularisation scale $\mu$. In this work we will take $\mu=1000~{}{\rm MeV}$ consistently with previous approaches Montana:2017kjw . Although these two expression for the loop function, (14) and (15), give a rather different result at low (sub-threshold) and high energies, they can be rather similar in the threshold region of the corresponding channels. The value of the subtraction constant that produces the same loop function at threshold as that obtained with a cut-off $\Lambda$ is determined by
@@
Some mesons in the considered channels have a large width $(\Gamma_{\rho}=149.4~{}{\rm MeV}$ and $\Gamma_{K^{*}}=50.5~{}{\rm MeV}$), but $G_{l}$ in Eq. (15) does not take into account this fact. Therefore, in order to implement the meson width we convolute this function with the mass distribution of the particle. This method has been used in Montana:2017kjw ; Oset:2010tof and the resulting loop function for these channels becomes:
@@
Then, the obtained scattering amplitude ($T$) can be analytically continued to the complex plane of $s$. The dynamically generated resonances appear as poles of the amplitude $T_{ij}$ in the so-called second Riemann sheet, obtained by using the following rotation of the loop function Montana:2017kjw :
@@
from which we can obtain the coupling constants, $g_{i}$, of the pole/resonance for all channels. In addition, we calculate the compositeness defined as:
$\chi_{i}=\Bigl{\lvert}g_{i}^{2}\frac{\partial G_{i}(\sqrt{s})}{\partial\sqrt{s}}\Bigr{\rvert}_{\sqrt{s}=\sqrt{s_{p}}},$ (22)
which approximately gives the contribution of the $i$-th channel meson-baryon component in a given resonance.

### v2

@@ Abstract
We study the possibility that four $\Xi$ resonances $(\Xi(1620)$, $\Xi(1690)$, $\Xi(1820)$, $\Xi(1950))$ could correspond to pentaquark states, in the form of meson-baryon bound systems. We also explore the possible existence of doubly strange pentaquarks with hidden charm $(P_{css})$ and find two candidates structured in a similar form, at energies of $4493~{}{\rm MeV}$ and $4630~{}{\rm MeV}$. The meson-baryon interaction is built from t-channel meson exchange processes which are evaluated using effective Lagrangians. Moreover we analyse the $\Xi_{b}\to\Xi~{}J/\psi~{}\phi$ decay process, which permits exploring the existence of the heavy double strange pentaquark, as well as other exotic hadrons, in the three different two-body invariant mass spectra of the emitted particles. In the $J/\psi\phi$ mass spectrum, we analyse the nature of the $X(4140)$ and $X(4160)$ resonances. In the $J/\psi\Xi$ invariant mass spectrum, we study the signal produced by the double strange pentaquark, where we conclude that it has a good chance to be detected in this reaction if its mass is around $4580-4680~{}{\rm MeV}$. Finally, in the $\phi\Xi$ spectrum we study the likelihood to detect the $\Xi(2500)$ state.
@@ I Introduction
One example of the success of this type of models is provided by the $\Lambda(1405)$ resonance, for which the quark models systematically over-predicted its mass. Instead, dedicated studies of the meson-baryon interaction obtained from chiral effective Lagrangians predict the $\Lambda(1405)$ to be a $\bar{K}N$ molecule [1, 2, 3], what actually allows to interpret the $\Lambda(1405)$ as the first observed pentaquark. Furthermore, the chiral models with unitarization in coupled channels have predicted its double-pole nature [4, 5], which can be seen from comparing different experimental line shapes [6], and which is now commonly accepted in the field and appears in the PDG  [7]. After the successful interpretation of the $\Lambda(1405)$, many groups tried to extend this kind of models to other spin, isospin and flavour sectors, finding a more natural way to explain different states, like the $f_{0}(500)$ and the $a_{0}(980)$, which are described as meson-meson quasi-bound states [8, 9, 10].
@@
More recent experimental results, like the discovery at LHCb [11, 12] of four excited resonances of the nucleon ($P_{c}(4312)$, $P_{c}(4380)$, $P_{c}(4440)$ and $P_{c}(4457))$, seen in the invariant mass distribution of $J/\psi p$ pairs from the $\Lambda_{b}$ decay, and the more recent report of LHCb [13] which, analysing the $\Xi_{b}$ decay, finds evidence of the existence of a pentaquark with strangeness $(P_{cs}(4459))$ on the invariant mass distribution of $J/\psi\Lambda$ pairs, clearly establish the need for including a $c\bar{c}$ pair excitation in order to reproduce the high masses of these states.
@@
Baryons with hidden charm, structured as a meson-baryon molecule, had been predicted in some earlier works [14, 15, 16, 17, 18], prior to their discovery, but the existence of a hidden charm pentaquark with double strangeness was not addressed until recently. Adopting a one-boson-exchange model to derive effective interaction potentials among the hadrons, the work of [19] found a pair of slightly bound meson-baryon molecules with double strangeness, coupling strongly to $\Xi_{c}^{\prime}\bar{D}_{s}^{*}$ and $\Xi_{c}^{*}\bar{D}_{s}^{*}$ states respectively, albeit with the use of somewhat hard and unrealistic cut-off values of around 2 GeV. In contrast, the works based on a t-channel vector-meson exchange interaction between mesons and baryons in the hidden charm sector did not find double strangeness pentaquark molecules [14]. However, a recent study pointed out that, in spite of the fact that these models predict weakly attractive or even repulsive meson-baryon interactions, a strong coupled-channel effect does indeed produce enough attraction to generate bound pentaquarks [20, 21] with strangeness $S=-2$, $P_{css}$. The prediction of such pentaquarks has been corroborated in Ref. [22], where a similar t-channel vector-meson interaction was employed, but deriving the baryon-baryon-vector meson vertices from the spin and flavour wave functions rather than from SU(4) arguments as in [20]. We note that $P_{css}$ pentaquarks have also been predicted within quark-model approaches [23, 24] or employing sum rule techniques [25, 26].
@@
In the recent work of Ref. [27] the authors have proposed the $\Xi_{b}^{0}\to\eta\eta_{c}\Xi^{0}$ and $\Omega_{b}^{-}\to K^{-}\eta_{c}\Xi^{0}$ decay processes to look for the $P_{css}$ with $I(J^{P})=\frac{1}{2}(\frac{1}{2}^{-})$ generated via the pseudoscalar-baryon interaction. The enhanced intensity of the $\Omega_{b}^{-}$ decay process establishes it as having a better chance for observation according to their study, but it is also much more difficult experimentally.
@@
This paper has a twofold purpose. On the one hand, we present the details of the formalism that produces dynamically generated $P_{css}$ states, which were omitted in a previously published short letter [20], while showing at the same time the prediction of the model for the lighter energy range of the $S=-2$ sector (including some other resonances generation). This sector has experienced a renewed interest due to recent data from the Belle experiment [28], studied, for example, in [29, 30], and even more recently from the femtoscopy analysis of the ALICE collaboration [31], which are analyzed in particular in Ref. [32].
@@
On the other hand, we propose a decay process, $\Xi_{b}\to\Xi~{}J/\psi~{}\phi$, in which the predicted pentaquark can be seen in the $J/\psi\Xi$ spectrum. As we will discuss in section IV.1, this reaction is interesting because it allows to detect and study some other exotic hadrons too, namely a family of $X$ resonances from $X(4140)$ till $X(4700)$ and $\Xi(2500)$.
@@ II.1 Meson-Baryon interaction
The model of the meson-baryon interaction used in this work is based on the tree-level diagrams of Fig. 1. We will only consider the s-wave amplitude, for which the most important contribution comes from the t-channel term (Fig. 1(a)). The s- and u- channel terms (Fig. 1(b) and (c)) contribute mostly to the p-wave amplitude and they may have effects at higher energies. In Ref. [4] the contribution from these terms in the $S=-1$ sector were calculated, and were found to be around $20\%$ of that of the dominant t-channel around $200~{}{\rm MeV}$ above the threshold. We can expect that in the $S=-2$ sector studied here, these terms will contribute even less, as the intermediate baryon is more massive, and therefore they will be neglected.
Figure 1: Leading order tree-level diagrams contributing to the meson-baryon interaction: t-channel term (a), s-channel term (b), and u-channel term (c). Baryons and mesons are depicted by solid and dashed lines, respectively.
To describe the interaction we employ effective Lagrangians that describe the couplings of the vector meson to pseudoscalar mesons $(VPP)$ and baryons $(VBB)$, which are obtained using the hidden gauge formalism and assuming $SU(4)$ symmetry [15]:
@@
Using the $VPP$ and $VBB$ vertices in the t-exchange diagram of Fig. 1 (a) one can obtain the s-wave interaction kernel $V_{ij}$ [15] for the pseudoscalar meson baryon $(PB)$ interaction:
@@ II.2 The $I=1/2~{}S=-2$ sector
In the $I=1/2~{}S=-2$ sector we have $9$ possible pseudoscalar-meson baryon (PB) channels, namely $\pi\Xi(1456)$, $\bar{K}\Lambda(1611)$, $\bar{K}\Sigma(1689)$, $\eta\Xi(1866)$, $\eta^{\prime}\Xi(2276)$, $\eta_{c}\Xi(4302)$, $\bar{D}_{s}\Xi_{c}(4437)$, $\bar{D}_{s}\Xi_{c}^{\prime}(4545)$, $\bar{D}\Omega_{c}(4565)$, where the values in parenthesis denote their thresholds. As can be seen, the channels with hidden charm are about $2~{}{\rm GeV}$ more massive compared with the lighter channels. Therefore we can expect these two regions to be essentially independent, making only a rather small effect on each other. The values of the $C_{ij}$ coefficients for the nine channels are displayed in Table 1.
@@
$\eta^{\prime}\Xi$ 00 00 $-\frac{1}{\sqrt{2}}\kappa_{c}$ $-\frac{1}{\sqrt{6}}\kappa_{c}$ $\frac{1}{\sqrt{3}}\kappa_{c}$
@@
In the vector meson baryon (VB) sector, the allowed channels are $\rho\Xi(2089)$, $\bar{K}^{*}\Lambda(2010)$, $\bar{K}^{*}\Sigma(2087)$, $\omega\Xi(2101)$, $\phi\Xi(2338)$, $J/\psi\Xi(4415)$, $\bar{D}^{*}_{s}\Xi_{c}(4581)$, $\bar{D}^{*}_{s}\Xi_{c}^{\prime}(4689)$, $\bar{D}^{*}\Omega_{c}(4706)$, which can again be separated in two groups, one with the light channels and the other with the heavy ones. In this case we can obtain the $C_{ij}$ coefficients directly from those in Table 1 using the following correspondences,
$\pi\rightarrow\rho,~{}~{}K\rightarrow K^{\ast},~{}~{}\bar{K}\rightarrow\bar{K}^{\ast},~{}~{}D\rightarrow D^{\ast},~{}~{}\bar{D}\rightarrow\bar{D}^{\ast}$
$\eta_{c}\rightarrow J/\psi,\frac{1}{\sqrt{3}}\eta+\sqrt{\frac{2}{3}}\eta^{\prime}\rightarrow\omega,-\sqrt{\frac{2}{3}}\eta+\frac{1}{\sqrt{3}}\eta^{\prime}\rightarrow\phi\ ,$ (9)
@@
$\omega\Xi$ 00 00 00 00 00 $\frac{1}{\sqrt{2}}\kappa_{c}$
$\phi\Xi$ 00 00 $-\sqrt{\frac{3}{2}}\kappa_{c}$ $-\frac{1}{\sqrt{2}}\kappa_{c}$ 00
@@
Note that in the light $PB$ sector there are formally only $4$ coupled channels, due to the fact that all $C_{5j}$ coefficients with $j\leq 5$ are zero, while this is not the case for the light $VB$ sector, which remains with $5$ coupled channels.
@@ II.3 Dynamically generated resonances
The sought resonances are dynamically generated as poles of the scattering amplitude $T_{ij}$, which is unitarized via the on-shell Bethe-Salpeter equation in coupled channels. This corresponds to performing the resummation of loop diagrams to an infinite order (see Fig. 2),
@@
The masses $M_{l}$ and $m_{l}$ correspond to those of the baryon and meson of the $l$ channel, respectively, $P=p+k=(\sqrt{s},0)$ is the total four momenta in the c.m. frame, and $q$ denotes the four momentum in the intermediate loop. We factorize the $V$ and $T$ matrices on-shell out of the integrals in Eq. (10) [33], hence obtaining the following algebraic matrix expression for the scattering amplitude,
@@
Note that, in searching the poles of the VB amplitudes, one can consider only the transverse mode of the vector polarization vector, as discussed in Appendix B of Ref. [34]. Moreover, disregarding the small contribution of the $q_{i}q_{j}/M_{V}^{2}$ term of the virtual vector meson propagator, one can factorize the product $\vec{\epsilon}_{i}\vec{\epsilon}_{j}$ out from all terms in the Bethe-Salpeter equation.
@@ II.4 Loop function
were $q_{l}$ is the on-shell three-momentum of the meson in the loop, and we have introduced a subtraction constant, $a(\mu)$, at the regularisation scale $\mu$. In this work we will take $\mu=1000~{}{\rm MeV}$ consistently with previous approaches [33]. Although these two expression for the loop function, (13) and (14), give rather different results at low (sub-threshold) and high energies, they can be rather similar in the threshold region of the corresponding channels. The value of the subtraction constant that produces the same loop function at threshold as that obtained with a cut-off $\Lambda$ is determined from
@@
Some mesons in the considered channels have a large width $(\Gamma_{\rho}=149.4~{}{\rm MeV}$ and $\Gamma_{K^{*}}=50.5~{}{\rm MeV}$), and this is implemented by convoluting $G_{l}$ in Eq. (14) with the mass distribution of the particle. This method has been used in [33, 35] and the resulting loop function for the channels involving wide mesons becomes:
@@
Then, the obtained scattering amplitude ($T$) can be analytically continued to the complex plane of $s$. The dynamically generated resonances appear as poles of the amplitude $T_{ij}$ in the so-called second Riemann sheet, obtained by using the following rotation of the loop function [33]:
@@
from which we can obtain the coupling constants, $g_{i}$, of the pole/resonance for all channels.
In addition, we provide a measure of the content of the $i^{\rm th}$ channel meson-baryon component in a given resonance. Let us note that the “compositeness” of resonances, defined as
$X_{i}=\left.g_{i}^{2}\displaystyle\frac{\partial G_{i}(\sqrt{s})}{\partial\sqrt{s}}\right|_{\sqrt{s}=\sqrt{s_{p}}}$ (21)
in analogy to that of bound states, is a complex quantity that represents the integral of the $i^{\rm th}$ component of the wave-function squared, not the modulus squared (see Ref. [36] for a summary of the present understanding of the compositeness in hadrons). Therefore, strictly speaking, the former definition cannot be interpreted as a probability, and some prescriptions have been given in the literature to connect this complex quantity to a real one, $\chi_{i}$, which provides a measure of the strength of the $i^{\rm th}$ channel within the resonance wave-function. We have taken $\chi_{i}=\left|X_{i}\right|$, which is the interpretation derived in Ref. [37] and requires the resonance mass to be larger than the decaying channels. So, formally, our results for $\chi_{i}$ in Tables III, V, VI and VII should only have a meaningful probabilistic interpretation for channels below the quoted resonance mass. An alternative definition is to associate $\chi_{i}$ to the real pat of $X_{i}$, as adopted in Ref. [38], a prescription which ensures the fulfillment of the sum rule $\sum_{i}\chi_{i}+Z=1$, where $Z$ is the amount of the genuine (non-molecular) component of the resonance. Other prescriptions for real valued compositeness in resonances can be found in [39, 40].